# RGMP: Recurrent Geometric-prior Multimodal Policy for Generalizable Humanoid Robot Manipulation

### v1

@@ 4 Experiments
In this section, we evaluate the effectiveness and generalized ability of the proposed RGMP framework. Our experiments are designed to assess the performance of each core component, including the GSS and ARGN. We also compare our approach against state-of-the-art baselines in various manipulation tasks. The evaluation metrics, experimental setup, implementation details, and comprehensive results are detailed as follows.
@@ 4.1 Hardware Setup
We conduct experiments on two robotic platforms: a humanoid robot, with evaluations focused on the upper limb (please see Appendix D in supplements for details), and a desktop dual-arm robot, developed to test cross-embodied generalized ability. The desktop robot is equipped with an RGB camera and two 6-DoF arms for manipulation tasks.
@@ 4.2 Dataset and Evaluation Criteria
To validate the effectiveness of the RGMP, we collected 120 trajectories for the skill library. Each trajectory corresponds to an execution path associated with an RGB image captured prior to the robotic arm performing an action. Specifically, each RGB image is linked to a target action, and each trajectory represents the movement of the robotic arm from its initial position to the target spatial location:
where denote the joint space of the robotic arm, where each trajectory specifies the motion of the arm from its initial configuration to the target spatial location and pose of the end effector. In real-world evaluations, the model performance is assessed using two complementary metrics. The skill success rate, denoted as , is recorded when the robot correctly identifies and selects the appropriate skill for the task. Additionally, the execution accuracy quantifies the precision with which the robot executes the selected skill to retrieve the target object. Consequently, the final success rate is defined as the product of these two metrics:
The detailed criteria for ManiSkill2 manipulation tasks can be referred to in Appendix E in supplements.
@@ 4.3 Performance Comparison and Ablation study
To evaluate RGMP, we conducted real-world comparative experiments against ResNet50 (he2016deep), Transformer (vaswani2017attention), the first-place entry (gao2023two) in the Maniskill2 challenge, Octo (team2024octo), OpenVLA (kim2024openvla), RDT-1b (liu2024rdt), Dex-VLA (wen2025dexvla), and Diffusion Policy. Experiments involved random target object placement, with success defined as accurate instruction understanding, correct manipulation execution (object delivery to humans), and collision avoidance with surrounding objects. To assess generalization and cross-domain transferability, we deployed the trained model on a desktop dual-arm platform, using a low-data setup: 40 interaction samples for Fanta can grasping as the exclusive training data, with evaluation on three unseen categories (human hands, spray bottles, Coke cans) at random workspace positions. Tables 1 and 2 show RGMP outperforms baselines across tasks, with top , , for Fanta, Sprite, tissues, squashed Coke cans, human hands, validating its effectiveness on regular/irregular objects. As Table 1 demonstrates, our GSS yields a 15-25% accuracy improvement in skill selection compared to Qwen-vl. Ablation studies (Table 3) confirm that integrating GMM with ARGN enhances performance: for Diffusion Policy, GSS+GMM yields a 0.55 versus 0.49 for without GMM, while ARGN with GSS+GMM achieves a 0.69 in picking squashed Coke, demonstrating the effectiveness of GMM in refining predictions. Additionally, Table 4 validates the contributions of RoPE, Spatial Mixing Blocks (SMB), and Channel Mixing Blocks (CMB): their combined use yields the highest accuracy across all objects (0.98 for Fanta, 0.78 for Coke, 0.81 for spray, 0.90 for hands). Beyond three primitives (grasp/lift-up/pinch), we evaluate five non-grasp ManiSkill2 tasks, complex tasks like plugging chargers (pinch) and opening cabinets (grasp) are dynamically composed from our atomic primitives. As shown in Figure 6, our RGMP achieves the highest performance across all tasks, demonstrating its transferability and generalization capability. Furthermore, as shown in Table 5, RGMP achieves a score of 0.98 with 40 training samples, which is 5× fewer than the 200 samples required by DP.
@@ 5 Conclusion and Future Work
This work addresses semantic-spatial skill alignment and visuomotor overfitting in humanoid robotics via RGMP, an end-to-end framework integrating GSS and ARGN. By dynamically associating contextual skills and decomposing 6-DoF trajectories into probabilistically regularized Gaussian components, RGMP achieves 87% generalization success and 5× greater data efficiency than Diffusion Policy in human-robot interaction. Results show explicit neuro-symbolic coordination enables robust generalization across unseen objects/scenes, advancing collaboration with a scalable adaptive manipulation foundation. Future work will explore functional generalization: demonstrating one primary object function allows automatic inference of trajectories for others, eliminating exhaustive teaching and enhancing efficiency in dynamic environments.
@@ 6 Acknowledgement
This work is supported by National Natural Science Foundation of China under Grant (62361166629, 62225113, 623B2080), the Major Project of Science and Technology Innovation of Hubei Province (2024BCA003, 2025BEA002), the Innovative Research Group Project of Hubei Province under Grants 2024AFA017, and the Key Research Project of Wuhan City 2024060788020073. The supercomputing system at the Supercomputing Center and the Learning Algorithms & Soft Manipulation Laboratory of Wuhan University supported the numerical calculations and the robot platforms in this paper.
Appendix A Implementation Details of GSS
A.1 GSS Architecture Overview
The GSS framework is implemented as a modular pipeline with three core components: (1) a Visual-Language Interpretation Module, (2) a Semantic Segmentation Module, and (3) a Prompt-Based Skill Selector. These components work in tandem to realize the translation from user instructions and visual observations to skill selection, with a total inference latency of 105 ms on an NVIDIA 4090 GPU.
Visual-Language Interpretation Module:
This module leverages the Qwen-vl API (instead of a locally deployed model) for visual-language understanding. Prior to deployment, a critical step involves loading historical knowledge into Qwen-vl through structured prompt engineering. This process entails feeding the API a curated set of training examples, each consisting of: (1) an RGB image of an object with its bounding box, (2) the corresponding semantic segmentation result of the object, (3) the manually annotated shape category (e.g., “cylindrical”, “crushed”), and (4) the optimal manipulation skill. For instance, one example includes the RGB image of a Fanta can with its bounding box, the segmentation result highlighting the contours of can, the label “cylindrical shape”, and the associated skill “side Grasp”. Another example comprises the RGB image of a crushed cola can with its bounding box, the segmentation result showing the flattened structure, the label “crushed shape”, and the skill “Lift up”. A total of 20 such examples covering common object categories are used to prime the model, enabling it to learn the mapping between visual features, shape attributes, and manipulation skills. The input prompt template for the API after knowledge loading is structured as: “Instruction: Identify target object in image and output bounding box [x1, y1, x2, y2] ”. The API achieves 93.1% object localization accuracy on our custom dataset, with an average response time of 45 ms per query.
Semantic Segmentation Module:
After obtaining the bounding box coordinates from the Qwen-vl API, the corresponding object is cropped out and input into the YOLOv8-seg model for semantic segmentation. This model, pretrained on a large-scale dataset and fine-tuned on our custom dataset, extracts the shape information of the target object. It achieves a mean Intersection over Union (mIoU) of 97.6% on our custom dataset, ensuring accurate segmentation results for shape analysis.
Prompt-Based Skill Selector:
This component takes the bounding box coordinates and shape information of the target object as inputs and uses a specific prompt to make skill selections via the Qwen-vl API. The prompt is structured as: “You are a robot with three Skills: side Grasp, Lift up, and top Pinch. The image you are observing has a resolution of 640x480. Based on the observation, the Coordinates of the bounding box and the Shape information of the object, Choose the skill without collision.”
The selector utilizes the Qwen-vl API, which has learned from the prompt engineering examples, to map the input information to the appropriate skill. It resolves potential ambiguities by prioritizing skills that minimize collision risks based on the position of object and surrounding environment, with the API returning a confidence score for each candidate skill.
A.2 Skill Library
The skill library is defined based on various combinations of finger grip bending angles, enabling dexterous hands to execute different grasping techniques: side grasping, picking up, and pinching. The side grasping skill is applicable in scenarios where the robot arm can grasp objects from the side without any obstacles obstructing its path. The lifting up skill is used when an object must be lifted from above, typically in the presence of blocking obstacles. The pinching skill is intended for situations where the target object is small or thin, such as napkins, charging cables, and similar items. The skills within the library specifically define the grasping posture of the dexterous hand, without imposing strict constraints on the nature of the objects being grasped. The GSS selects the appropriate skill from the library based on both geometric-prior common sense and real-time RGB image observations, determining the feasible dexterous hand posture for the current scenario. Each skill is associated with distinct model weight parameters, and the action trajectory is dynamically generated in real-time by the RGMP.
Appendix B Details of ARGN
Particularly, the ARGN architecture comprises three principal components: a Stem layer for initial feature extraction, three hierarchical processing stages (each with two ARGN blocks), and a multi-scale feature fusion module integrated with GMM. This modular design enables ARGN to progressively refine visual features while maintaining computational efficiency.
For an input RGB image , the Stem layer performs spatial down-sampling and feature embedding to produce with a resolution of . This transformation is achieved through a series of convolutional operations, including a convolution, batch normalization (BN), Squared ReLU (SReLU) activation, and max-pooling. The Stem layer serves two key purposes: it reduces the spatial dimensionality to alleviate computational burdens in subsequent stages, and it preserves essential low-frequency visual information (e.g., object contours and global scene structure) that is vital for robust action generation. The three hierarchical processing stages (Stage 1 to Stage 3) each consist of two ARGN blocks, which are designed to iteratively enhance feature representations through spatial and channel-wise modeling. Each ARGN block comprises a Spatial Mixing Block and a Channel Mixing Block, working in tandem to capture both local spatial correlations and cross-channel dependencies. Through recursive operations within these blocks, ARGN builds a dynamic memory of visual patterns, allowing it to adaptively prioritize task-relevant regions (e.g., target objects or end-effector positions) in unseen environments. In order to reconstruct the spatial relationship, we encode the image patches with the Rotary Position Embedding (RoPE), which establishes an implicit association between each observed image patch and the final executed action by encoding positional information through geometric transformations. Formally, RoPE is defined as:
where is a pixel from the initial feature map , and denote the spatial indices (height and width) of , and represents the two-dimensional position code. This position code is formulated as:
Here, the frequency parameter governs rotation speeds across channel dimensions, with denoting the channel index. The base constant 10000 and exponential decay term jointly establish a multi-scale spectral distribution mechanism: low-frequency components (corresponding to small ) prioritize modeling global structural relationships, while high-frequency components (larger ) capture fine-grained local positional variations. This design allows RoPE to inherently encode relative spatial offsets without relying on learnable position-specific parameters, thereby reducing overfitting risks and enhancing generalization to novel scenes. By combining the feature initialization of Stem layer, iterative refinement of hierarchical ARGN blocks, and positional encoding of RoPE, ARGN effectively models the spatial dependencies between visual inputs and robotic actions, laying the foundation for accurate and adaptive action generation in complex environments.
Appendix C Detailed Derivation of GMM for Action Refinement
In our framework, the Gaussian Mixture Model (GMM) is trained on demonstration data to model the distribution of ground-truth joint angles, featuring 6 components that correspond to the 6 robotic joints. The model parameters are estimated through the Expectation-Maximization (EM) algorithm, which iteratively maximizes the log-likelihood of the observed data. The key steps are elaborated as follows:
E-Step (Expectation).
For each data point (i.e., a sample of ground-truth joint angles) and each component , the posterior probability (responsibility) that belongs to component is calculated as:
where represents a latent variable indicating whether is generated by component .
M-Step (Maximization).
The parameters are updated to maximize the expected log-likelihood, which is computed using the responsibilities obtained from the E-step. The weight of component is determined as the average responsibility of that component across all data points:
with denoting the number of demonstration samples. The mean of component is the weighted average of the data points, with the weights being :
| Methods | Push Chair | Move Bucket | Plug Charger | Cabinet Door | Cabinet Drawer | Average Score |
| RDT-1b | 0.04 0.01 | 0.01 0.00 | 0.01 0.00 | 0.13 0.03 | 0.10 0.02 | 0.06 0.01 |
| Dex-VLA | 0.02 0.00 | 0.03 0.01 | 0.01 0.01 | 0.18 0.02 | 0.08 0.02 | 0.06 0.01 |
| Maniskill2-1st | 0.08 0.02 | 0.04 0.01 | 0.03 0.01 | 0.24 0.03 | 0.12 0.02 | 0.10 0.02 |
| Octo | 0.05 0.01 | 0.02 0.01 | 0.03 0.01 | 0.20 0.03 | 0.10 0.01 | 0.08 0.01 |
| OpenVLA | 0.06 0.01 | 0.04 0.01 | 0.04 0.02 | 0.24 0.03 | 0.15 0.01 | 0.10 0.02 |
| Diffusion Policy | 0.04 0.01 | 0.06 0.02 | 0.06 0.02 | 0.15 0.01 | 0.17 0.02 | 0.10 0.01 |
| RGMP(ours) | 0.14 0.02 | 0.08 0.01 | 0.07 0.01 | 0.26 0.03 | 0.20 0.03 | 0.15 0.02 |
The covariance of component is the weighted covariance of the data points around :
The E-step and M-step are repeated until the log-likelihood converges, i.e., the change in drops below a predefined threshold (e.g., ). The GMM regression process refines the initial action (generated by the GSA model) by aligning it with the most consistent Gaussian component. This involves deriving the conditional distribution of joint angles given , along with the aggregated mean and covariance.
Conditional Distribution for Single Gaussian Component.
For the -th Gaussian component, the joint distribution of and is modeled as a multivariate normal distribution , where and are sub-means, and are sub-covariances. According to the properties of multivariate Gaussians, the conditional distribution for component is also a normal distribution:
In our setup, due to the high correlation between and (both representing joint angle predictions), we simplify by assuming is a noisy observation of . This leads to the conditional mean and covariance (as utilized in the main text).
Aggregated Mean and Covariance Across Components.
The overall conditional distribution is a weighted sum of the component-wise distributions, with weights (estimated from the EM algorithm). The aggregated mean:
where is re-normalized based on the consistency of with (measured via Mahalanobis distance). The aggregated covariance, which accounts for component weights:
where squaring ensures that components with lower weights contribute less to the overall uncertainty. The Mahalanobis distance between and is preferred over the Euclidean distance for two main reasons: 1) it accounts for the covariance structure, and 2) it normalizes for units. Unlike the Euclidean distance, which treats all dimensions equally, the Mahalanobis distance incorporates the covariance matrix , scaling distances by the variability of each joint angle. This is crucial for robotic joints, as some angles (e.g., those controlling fine movements) may have smaller variances and require stricter consistency checks. Additionally, joint angles can have different units or ranges (e.g., radians for rotation vs. meters for translation). The Mahalanobis distance normalizes these differences, enabling a meaningful comparison across dimensions. Formally, the distance quantifies how “atypical” is relative to the distribution of the -th component. A smaller indicates higher consistency, making the component a better candidate for refining the action prediction.
Appendix D Detailed Hardware Setup
The upper limb hardware comprises four components: visual input, voice input/output, behavioral decision-making, and action execution. The visual input consists of a self-centered RGB camera installed on the head of robot, while the voice input/output hardware employs the iFLYTEK S0Y22F omnidirectional microphone and speaker, also mounted on the chest of robot. The behavioral decision-making hardware utilizes an NVIDIA Orin chip installed in the head of robot. The motion execution hardware includes two 6-degree-of-freedom (6-DoF) robotic arms and two 6-DoF dexterous hands. The entire robot is self-powered and communicates via a local area network. The robot can be started and shut down without the external power or communication cables (as shown in Figure A3).
Appendix E ManiSkill2 Criteria
| Symbol | Type | Description |
| Training and Interaction | ||
| Scalar | Total training epochs for RGMP parameter optimization | |
| Integer | Maximum human-robot conversation rounds | |
| Integer | Capacity of human demonstration collection | |
| Inputs and Perception | ||
| Text | Transcribed human speech instruction | |
| Image | Ego-centric RGB camera frame of workspace | |
| Vector | 6-DoF robotic arm joint angles (radians) | |
| Dataset and Models | ||
| Collection | Demonstration tuples for training | |
| Model | Fine-tuned Qwen-vl vision-language model | |
| Policy | Pre-trained RGMP model for skill | |
| Features and Mechanisms | ||
| Tensor | Initial visual feature map from stem network | |
| Scalar | Adaptive decay factor in recursive memory | |
| Tensors | Key/value matrices for spatial attention mixing | |
| – | Tensors | Multi-scale feature maps from ARGN stages |
| Action Predictions | ||
| Vector | Initial action from ARGN network | |
| Vector | Human-demonstrated joint angles (ground truth) | |
| Scalar | MSE loss between predictions and ground truth | |
| Gaussian Mixture Model | ||
| Integer | Number of Gaussian components (6, matching arm DoFs) | |
| Scalar/Vector/Matrix | Prior, mean, and covariance of -th component | |
| Set | GMM parameters | |
| Vector | -th ground-truth joint angle sample | |
| Scalar | Responsibility of to component | |
| Integer | Total demonstration samples | |
| Scalar | Mahalanobis distance ( to ) | |
| Evaluation Metrics | ||
| Scalar | Skill selection accuracy (0-1) | |
| Scalar | Task execution accuracy (0-1) | |
| Scalar | Overall success rate () | |
Open Cabinet Drawer
A single-arm mobile robot opens a specified cabinet drawer (with randomized joint friction or damping); success requires the target drawer to be opened to 90% of its max range and remain static. The training set includes 25 cabinets (multiple drawers) with 300 trajectories per drawer; evaluation uses 10 unseen test cabinets in 2 stages (250 trajectories each: 125 from 5 unseen cabinets, 125 from training set). The target drawer is specified by its initial center of mass.
Open Cabinet Door
A single-arm mobile robot opens a specified cabinet door (with randomized joint friction or damping); success requires the target door to be opened to 90% of its max range and remain static. The training set includes 42 cabinets (multiple doors) with 300 trajectories per door; evaluation uses 10 unseen test cabinets following the same protocol as Open Cabinet Drawer. The target door is specified by a segmentation mask.
Push Chair
A dual-arm mobile robot pushes a swivel chair to a red hemisphere target without tipping it (with randomized joint parameters); success requires the chair to be within 15cm of the target, remain static, and stay upright. The training set includes 26 chairs with 300 trajectories per chair; evaluation uses 10 unseen chairs following the same protocol as Open Cabinet Drawer. No task-specific observations.
Move Bucket
A dual-arm mobile robot moves a ball-containing bucket and lifts it onto a platform; success requires the bucket to be upright on/above the platform, remain static, and retain the ball. The training set includes 29 buckets with 300 trajectories per bucket; evaluation uses 10 unseen buckets following the same protocol as Open Cabinet Drawer. No task-specific observations.
Plug Charger
A robot plugs a charger into a wall receptacle; success requires full insertion. The training set has 1,000 successful trajectories; evaluation has 2 stages (100 episodes each) with varied robot joint positions and charger poses. No task-specific observations.

### v2

@@ 4 Experiments
We evaluate the effectiveness and generalization of the RGMP framework by assessing its core components (GSS and ARGN) and benchmarking it against state-of-the-art methods. The evaluation metrics, experimental setup, implementation details, and comprehensive results are as follows:
@@ 4.1 Hardware Setup
We conduct experiments on two robotic platforms: a humanoid robot, with evaluations focused on the upper limb (please see Appendix D for details), and a desktop dual-arm robot, designed to test cross-embodiment generalization capability. The desktop robot is equipped with an RGB camera and two 6-DoF arms for manipulation tasks.
@@ 4.2 Dataset and Evaluation Criteria
To validate the effectiveness of the RGMP, we collected 120 trajectories for the skill library. Each trajectory corresponds to an execution path associated with an RGB image captured prior to the robotic arm performing an action:
where denotes the joint space of the robotic arm, each trajectory specifies the motion of the arm from its initial configuration to the target spatial location and end-effector pose. In real-world evaluations, the model performance is assessed using two complementary metrics. The skill success rate, denoted as , is recorded when the robot correctly identifies and selects the appropriate skill for the task. Additionally, the execution accuracy quantifies the precision with which the robot executes the selected skill to retrieve the target object. Consequently, the final success rate is defined as the product of these two metrics:
The detailed criteria for ManiSkill2 manipulation tasks can be referred to in Appendix E in the code link.
@@ 4.3 Performance Comparison and Ablation Study
To evaluate RGMP, we conducted real-world comparative experiments against ResNet50 (he2016deep), Transformer (c:22), the first-place entry (gao2023two) in the ManiSkill2 challenge, Octo (team2024octo), OpenVLA (kim2024openvla), RDT-1b (liu2024rdt), Dex-VLA (wen2025dexvla), and Diffusion Policy. Experiments involved random target object placement, with success defined as accurate instruction understanding, correct manipulation execution (object delivery to humans), and collision avoidance. To assess generalization and cross-domain transferability, we deployed the trained model on a desktop dual-arm platform, using a low-data setup: 40 interaction samples for Fanta can grasping as the exclusive training data, with evaluation on three unseen categories (human hands, spray bottles, Coke cans) at random workspace positions. Tables 1 and 2 show RGMP outperforms baselines across tasks, with top , , and for Fanta cans, Sprite cans, tissue papers, deformed Coke cans, and human hands, validating its effectiveness on regular/irregular objects. As Table 1 demonstrates, our GSS yields a 15-25% accuracy improvement in skill selection compared to Qwen-VL. Ablation studies (Table 3) confirm that integrating GMM with ARGN enhances performance: for Diffusion Policy, GSS+GMM yields a 0.55 versus 0.49 without GMM, while ARGN with GSS+GMM achieves a 0.69 in picking squashed Coke, demonstrating the effectiveness of GMM in refining predictions. Additionally, Table 4 validates the contributions of RoPE, Spatial Mixing Blocks (SMB), and Channel Mixing Blocks (CMB): their combined use yields the highest accuracy across all objects (0.98 for Fanta, 0.78 for Coke, 0.81 for spray, 0.90 for hands). Beyond three primitives (grasp/lift-up/pinch), we evaluate five non-grasp ManiSkill2 tasks, complex tasks like plugging chargers (pinch) and opening cabinets (grasp) are dynamically composed from our atomic primitives. As shown in Figure 6, our RGMP achieves the highest performance across all tasks, demonstrating its transferability and generalization capability. Furthermore, as shown in Table 5, RGMP achieves a score of 0.98 with 40 training samples, using 5 fewer samples than the 200 required by DP.
@@ 5 Conclusion and Future Work
This work addresses semantic-spatial skill alignment and visuomotor overfitting in humanoid robotics via our RGMP, an end-to-end framework integrating GSS and ARGN. By dynamically associating contextual skills and decomposing 6-DoF trajectories into probabilistically regularized Gaussian components, RGMP achieves 87% generalization success and 5 greater data efficiency than Diffusion Policy in human-robot interaction. Results show explicit neuro-symbolic coordination enables robust generalization across unseen objects/scenes, advancing collaboration with a scalable adaptive manipulation foundation. Future work will explore functional generalization: demonstrating one primary object function allows automatic inference of trajectories for others, eliminating exhaustive teaching and enhancing efficiency in dynamic environments.
@@ Acknowledgements
This work is supported by National Natural Science Foundation of China under Grants (62361166629, 62225113, 623B2080), the Major Project of Science and Technology Innovation of Hubei Province (2024BCA003, 2025BEA002), the Innovative Research Group Project of Hubei Province under Grant 2024AFA017, and the Key Research Project of Wuhan City 2024060788020073. The supercomputing system at the Supercomputing Center and the Learning Algorithms & Soft Manipulation Laboratory of Wuhan University supported the numerical calculations and the robot platforms in this paper.